# REFRAG: Rethinking RAG based Decoding

### v1

@@ 2 Model Architecture
Latency and throughput improvement. We evaluate three metrics: TTFT, the latency to generate the first token; TTIT, the time to generate each subsequent token; and Throughput, the number of tokens generated per unit time. Theoretical analysis (section˜9) shows that for short context lengths, our method achieves up to $k\times$ acceleration in TTFT and throughput. For longer context length, acceleration reaches up to $k^{2}\times$ for both metrics. Empirically, as shown in figure˜2, with a context length of $16384$ (mid-to-long context), REFRAG with $k=16$ achieves $16.53\times$ TTFT acceleration with cache and $8.59\times$ without cache [^1] REFRAG without cache means that we recompute the chunk embedding for the context and take this latency into account., both surpassing CEPE ($2.01\times$ and $1.04\times$, respectively), while achieving 9.3% performance (measured by perplexity) compared to CEPE (table˜1). We achieve up to $6.78\times$ throughput acceleration compared to LLaMA, significantly outperforming CEPE. With $k=32$, TTFT acceleration reaches $32.99\times$ compared to LLaMA ($3.75\times$ compared to CEPE) while maintaining similar performance to CEPE (see figure˜8 and table˜2). More detailed discussion on empirical evaluation is in section˜9.
@@ 4 Experimental Results
With a compression rate of $16$, we achieve a $9.3\%$ average perplexity improvement over CEPE across four datasets [^2] Percentage calculated as $\frac{\textsc{LLaMA-No Context}-\text{Perplexity to inspect}}{\textsc{LLaMA-No Context}-\min(\textsc{LLaMA-Full Context},\textsc{LLaMA-32K})}$. Meanwhile, our method is $16.53\times$ faster than LLaMA in TTFT and $2.01\times$ faster than CEPE (section˜10.4). At a compression rate of $32$, our perplexity matches CEPE, while TTFT acceleration increases to $30.85\times$ over LLaMA and $3.75\times$ over CEPE.
@@
Table 1: Perplexity on output tokens $x_{s+1:s+o}$ given context tokens $x_{1:s}$ for different models. We use $s=2048$ and $o\in\{512,1024,2048\}$ here. Bolding are based on comparing baselines excluding LLaMA-Full Context and LLaMA-32K since they are expected to be the best (ideally). The lower the better ($\mathbf{\downarrow}$).
@@
Table 2: Perplexity on output tokens $x_{s+1:s+o}$ given different length of context. We use $s\in\{4096,8192,16384\}$ and $o=2048$ here. Bolding are based on comparing baselines excluding LLaMA-Full Context and LLaMA-32K since they are expected to be the best (ideally). The lower the better ($\mathbf{\downarrow}$).
@@
Figure 3: Perplexity on $x_{s+1:s+o}$ under varying compression rates by selectively compressing different percentages of chunks. We compare three selection methods: RL (trained policy), Perplexity-desc (heuristic: lower perplexity), Perplexity-asc (heuristic: higher perplexity), and Random (random selection).
@@ 6 Related Works
Recent research has extensively investigated novel model architectures to improve retrieval-augmented generation. Guu et al. (2020) introduced pre-training for retrieval-augmented masked language models. Building on this, Borgeaud et al. (2022) proposed a new architecture and pre-training paradigm for generative LLMs, leveraging cross-attention and end-to-end pre-training with retrieval from a trillion-token data store, achieving strong performance. Subsequent work by Shi et al. (2024) and Lin et al. (2024) focused on fine-tuning existing LLMs by prepending retrieved passages to prompts and employing ensemble methods for response generation. Additionally, Izacard & Grave (2021) introduced fusion-in-decoder, which uses an encoder to process each passage in parallel and concatenates the hidden states for generation via a decoder. This approach accelerates attention computation by removing cross-document attention, but does not apply compression in the decoder, which could further reduce latency.
@@
Recent research has investigated various strategies to reduce memory usage and accelerate latency in long-context generation for LLMs. Choromanski et al. (2021) introduced compressed attention, reducing attention complexity from quadratic to linear; however, this method does not address memory requirements. It is complementary to our approach and can be integrated to further improve latency. StreamingLLM(Xiao et al., 2024) proposed attention sinks to decrease KV cache memory for long-context generation, though this does not reduce latency during the pre-filling stage. CEPE (Yen et al., 2024) employs cross-attention to token embeddings from context tokens, reducing both KV cache memory and attention computations. However, CEPE is limited to prefix context applications, as it disrupts the causal structure of the context, making it unsuitable for tasks such as multi-turn RAG or summarization. Additionally, CEPE does not utilize token compression, resulting in similar or even increased decoding latency.

### v2

@@ 2 Model Architecture
Latency and throughput improvement. We evaluate three metrics: TTFT, the latency to generate the first token; TTIT, the time to generate each subsequent token; and Throughput, the number of tokens generated per unit time. Theoretical analysis (section˜9) shows that for short context lengths, our method achieves up to $k\times$ acceleration in TTFT and throughput. For longer context length, acceleration reaches up to $k^{2}\times$ for both metrics. Empirically, as shown in figure˜2, with a context length of $16384$ (mid-to-long context), REFRAG with $k=16$ achieves $16.53\times$ TTFT acceleration with cache and $8.59\times$ without cache [^1] REFRAG without cache means that we recompute the chunk embedding for the context and take this latency into account., both surpassing CEPE ($2.01\times$ and $1.04\times$, respectively), while achieving 9.3% performance (measured by log-perplexity) compared to CEPE (table˜1). We achieve up to $6.78\times$ throughput acceleration compared to LLaMA, significantly outperforming CEPE. With $k=32$, TTFT acceleration reaches $32.99\times$ compared to LLaMA ($3.75\times$ compared to CEPE) while maintaining similar performance to CEPE (see figure˜8 and table˜2). More detailed discussion on empirical evaluation is in section˜9.
@@ 4 Experimental Results
With a compression rate of $16$, we achieve a $9.3\%$ average log-perplexity improvement over CEPE across four datasets [^2] Percentage calculated as $\frac{\textsc{LLaMA-No Context}-\text{Log-perplexity to inspect}}{\textsc{LLaMA-No Context}-\min(\textsc{LLaMA-Full Context},\textsc{LLaMA-32K})}$. Meanwhile, our method is $16.53\times$ faster than LLaMA in TTFT and $2.01\times$ faster than CEPE (section˜10.4). At a compression rate of $32$, our log-perplexity matches CEPE, while TTFT acceleration increases to $30.85\times$ over LLaMA and $3.75\times$ over CEPE.
@@
Table 1: Log-Perplexity on output tokens $x_{s+1:s+o}$ given context tokens $x_{1:s}$ for different models. We use $s=2048$ and $o\in\{512,1024,2048\}$ here. Bolding are based on comparing baselines excluding LLaMA-Full Context and LLaMA-32K since they are expected to be the best (ideally). The lower the better ($\mathbf{\downarrow}$).
@@
Table 2: Log-Perplexity on output tokens $x_{s+1:s+o}$ given different length of context. We use $s\in\{4096,8192,16384\}$ and $o=2048$ here. Bolding are based on comparing baselines excluding LLaMA-Full Context and LLaMA-32K since they are expected to be the best (ideally). The lower the better ($\mathbf{\downarrow}$).
@@
Figure 3: Log-Perplexity on $x_{s+1:s+o}$ under varying compression rates by selectively compressing different percentages of chunks. We compare three selection methods: RL (trained policy), Perplexity-desc (heuristic: lower perplexity), Perplexity-asc (heuristic: higher perplexity), and Random (random selection).
@@ 6 Related Works
Recent research has extensively investigated novel model architectures to improve retrieval-augmented generation. Guu et al. (2020) introduced pre-training for retrieval-augmented masked language models. Building on this, Borgeaud et al. (2022) proposed a new architecture and pre-training paradigm for generative LLMs, leveraging cross-attention and end-to-end pre-training with retrieval from a trillion-token data store, achieving strong performance. Subsequent work by Shi et al. (2024) and Lin et al. (2024) focused on fine-tuning existing LLMs by prepending retrieved passages to prompts and employing ensemble methods for response generation. Additionally, Izacard and Grave (2021) introduced fusion-in-decoder, which uses an encoder to process each passage in parallel and concatenates the hidden states for generation via a decoder. This approach accelerates attention computation by removing cross-document attention, but does not apply compression in the decoder, which could further reduce latency.
@@
Recent research has investigated various strategies to reduce memory usage and accelerate latency in long-context generation for LLMs. Choromanski et al. (2021) introduced compressed attention, reducing attention complexity from quadratic to linear; however, this method does not address memory requirements. It is complementary to our approach and can be integrated to further improve latency. StreamingLLM(Xiao et al., 2024) proposed attention sinks to decrease KV cache memory for long-context generation, though this does not reduce latency during the pre-filling stage. CEPE (Yen et al., 2024) employs cross-attention to token embeddings from context tokens, reducing both KV cache memory and attention computations. However, CEPE is limited to prefix context applications, as it disrupts the causal structure of the context, making it unsuitable for tasks such as multi-turn RAG or summarization. Additionally, CEPE does not utilize token compression, resulting in similar or even increased decoding latency. Concurrently with our work, Dai et al. (2025) proposed PCC, an embedding-based memory mechanism that summarizes past context into compact vectors, enabling retrieval of salient information during subsequent processing. Like CEPE, PCC is limited to prefix context applications and does not support arbitrary folding or expansion of contexts at any position. Interestingly, Kuratov et al. (2025) investigated the capacity of LLMs to encode long contexts into a single embedding, demonstrating minimal information loss for sequences up to 1500 tokens. Their work examines the extent to which information can be compressed into a single embedding, offering a complementary perspective to REFRAG, which is designed for decoding from multiple compact embeddings within the standard decoder architecture.